# Toward quantum-safe scalable networks: an open, standards-aware key management framework

### v1

@@ Abstract
With the advent of quantum computing, the increasing threats to security poses a great challenge to communication networks. Recent innovations in this field resulted in promising technologies such as Quantum Key Distribution (QKD), which enables the generation of unconditionally secure keys, establishing secure communications between remote nodes. Additionally, QKD networks enable the interconnection of multi-node architectures, extending the point-to-point nature of QKD. However, due to the limitations of the current state of technology, the scalability of QKD networks remains a challenge toward feasible implementations. When it comes to long-distance implementations, trusted relay nodes partially solve the distance issue through the forwarding of the distributed keys, allowing applications that do not have a direct QKD link to securely share key material. Even though the relay procedure itself has been extensively studied, the establishment of the relaying node path still lacks a solution. This paper proposes an innovative network architecture that solves the challenges of Key Management System (KMS) identification, relay path discovery, and scalability of QKD networks by integrating Software-Defined Networking (SDN) principles, and establishing high-level virtual KMSs (vKMS) in each node and creating a new entity called the Quantum Security Controller (QuSeC). The vKMS serves the end-user key requests, managing the multiple KMSs within the node and abstracting the user from discovering the correct KMS. Additionally, based on the high-level view of the network topology and status, the QuSeC serves the path discovery requests from vKMSs, computing the end-to-end (E2E) relay path and applying security policies. The paper also provides a security analysis of the proposal, identifying the security levels of the architecture and analyzing the core networking security properties.
Index Terms:
QKD, Trusted Relay, SDN, network security, quantum communications.
@@ I Introduction
As quantum technologies evolve, Quantum Key Distribution (QKD) stands as a leading technology for securing communications with unconditional security [1]. Unlike classical cryptographic mechanisms that rely on computational assumptions, QKD is based on quantum mechanics to generate cryptographic keys with information-theoretic security.
While the typical application scenario of QKD is point-to-point by nature, QKD networks interconnect multiple nodes over broader areas, enabling key delivery beyond direct optical links. However, the scalability of such networks remains a challenge, especially as the number of users, links, loops, and relay hops increase.
@@
Figure 1 illustrates two distinct QKD network topologies. The linear topology (top) represents most of the state-of-the-art trusted relay deployments, built pairs of Simple QKD nodes. This approach is applicable to static end-to-end (E2E) communications in which the relay path remains constant [2] [3]. However, the partial mesh topology (down) illustrates a more realistic scenario, where multiple QKD nodes co-located in the same site enable the provision of dynamic relay paths.
@@
  • Path Discovery and Relay Coordination: In nodes hosting multiple KMSs, each KMS must be able to identify and coordinate with the correct KMS instance within the same node to perform the relay of keys.
@@
These issues stem from the limited knowledge at individual KMSs, and by the limitations of current QKD standards, which support basic functions but lack full dynamic multi-hop capabilities. To address these limitations, this paper proposes a secure multi-hop key delivery framework based on a centralized, programmable architecture leveraging Software-Defined Networking (SDN) principles. This approach abstracts complexity from applications and KMSs, introducing a global coordination mechanism to resolve paths and manage relay processes. The proposed design enables a centralized discovery and resolution of relay paths and remains compatible with existing and evolving QKD standards. Additionally, it offers the possibility to integrate seamlessly different manufacturers and technologies.
The rest of the paper is organized as follows. Firstly, Section II summarizes the background and literature research of trusted-relay based QKD networks. Section III introduces the proposed system and describes the operational flow. Subsequently, Section IV provides a theoretical security analysis of the proposed system. Finally, Section V presents the main conclusions of the work.
@@ II Related work
A primary limitation of QKD is its restricted transmission distance, typically 100–150 kilometers, due to signal attenuation and the lack of practical quantum repeaters. To overcome these constraints and support the deployment of scalable and complex network topologies, QKD networks commonly rely on trusted relay nodes, which have been widely studied [4] [5].
@@
Many QKD deployments are based on linear topologies, where each trusted relay node hosts two QKD modules and associated KMSs, making the relay procedure relatively straightforward. For instance, [2] presents a linear QKD network connecting cities in Italy, Slovenia, and Croatia through three links and two relays. Similarly, the Beijing-Shanghai Backbone QKD Network [3] connects 32 nodes linearly across 2.000 kilometers. While such networks demonstrate the feasibility of large-scale deployments, their linear nature limits scalability and flexibility.
These limitations become more pronounced in practical topologies where nodes may host multiple KMSs and QKD links, In such configurations, the process of discovering the best KMSs and computing an optimal relay path becomes significantly more complex. To address this, several efforts explore distributed routing approaches for QKD networks. For example, [6] and [7] propose the use of modified versions of classical routing protocols such as OSPFv2. In these systems, each node independently computes shortest path trees and maintains a local routing table to determine relay paths.
Other works pursue centralized architectures for QKD network control and key relay. For example, [8] proposes a Key Management Agent per node and a centralized KMS coordinating all the links and networking functions.
In this context, SDN emerges as a promising approach. By decoupling control and data planes, SDN enables centralized control, programmability, dynamic path computation, and real-time reconfiguration, features well-suited for QKD networks with varying key rates, link availability, and security constraints.
Several recent efforts validate SDN integration in QKD infrastructures. For instance, [9] presents a network comprised of 28 QKD modules distributed across 9 nodes in Madrid. In addition, all the network is coordinated by an SDN controller through the SDN Agents present in the nodes. However, their relay nodes do not have KMS functionality for key retrieval, only extending the QKD reach.
Additionally, [10] proposes an SDN-managed QKD infrastructure prototype where E2E keys are delivered via key relay procedure orchestrated by the controller, which computes suitable key relay paths. However, the example used in the key relay process description is based on a simple linear chain and the procedure of the path computation is not described in detail, which limits its applicability to more complex topologies.
In line with these SDN-based approaches, the ETSI GS QKD 015 Control Interface for Software Defined Networks [11] specification stands as a key enabler for building interoperable and programmable QKD networks. This specification defines an interface and data models between the Software-Defined QKD nodes and the controller. It formalizes the architectural principles to integrate SDN control into QKD infrastructures to optimize the creation and usage of QKD keys.
@@ III-A Architecture
  • Local KMS: KMS instance associated to a physical QKD module with standard KMS functionalities: key storage, synchronization, management and exchange. It also handles both direct key retrieval and multi-hop key relaying, in accordance with instructions received from the QuSeC. From this point forward, these instances will be referred to as simply KMS.
  • QuSeC: It maintains a global view of the QKD network topology, including QKD links, node roles, and KMS associations. It responds to discovery requests from vKMS instances, computes optimal relay paths and enforces relay policies to all involved KMSs across the network. The relay path computation is performed based on SPF algorithm [12], assigning the weights of each link according to configurable criteria such as available key rate in each link, hop count, or physical distance, allowing for flexible adaptation to different optimization goals.
This architecture is agnostic to the underlying QKD link technology, allowing deployment over optical fiber, free-space, or even satellite QKD links, and multiple QKD implementation technologies including Discrete-Variable QKD (DV-QKD), Continuous-Variable (CV-QKD) or entanglement-based. From a scalability perspective, the design supports horizontal scalability, enabling additional QKD nodes to be integrated without compromising the network and KMS operation.
An example of the proposed architecture is depicted in Figure 2, illustrating a network comprising four nodes. In this scenario, three nodes operate as trusted relay nodes, each hosting multiple KMS instances (each connected to a different QKD module, potentially from different vendors) to interface with separate QKD links. The fourth node acts as a simple QKD node with a single QKD module.
@@ III-B Operational flow
The proposed architecture enables E2E QKD key delivery through a mechanism that integrates centralized computation of the relay path. In addition, the mechanism has been designed to be compliant with current specifications, including the ETSI GS QKD 014 [13] for key delivery, the ETSI GS QKD 020 [14] for interoperable KMSs in horizontal key relay procedures, which is currently under development, and the ETSI GS QKD 015 [11] for SDN integration. This section outlines the main operational steps from the application’s key request to the delivery of the final cryptographic key. These steps are illustrated in Figure 2, which includes the operations required for initiating and target applications (APP_A and APP_B) in Figure 2(a) and Figure 2(b) respectively.
(a) Steps of the proposed mechanism for an initiating application.
(b) Steps of the proposed mechanism for a target application.
Figure 2: Steps of the proposed mechanism for initiating and target applications.
@@ III-C Message Exchange/examples
This section presents two representative message exchange sequences to illustrate the operation of the proposed mechanism under different scenarios: one using a direct key delivery and the other involving a trusted relay-based relay path. These examples focus on the interaction between the applications, vKMS, KMSs, and the QuSeC during the key request and delivery process.
@@
Since no further hops are installed in the KMS, the relay process in completed. In addition, completion notifications are propagated back through the chain. Once KMS_1b receives the confirmation of successful relay completion, it responds to the initial Get Key request, returning K1 to the vKMS, which then delivers it to APP_A.
APP_B’s behavior mirrors the direct case: it sends a Get Key with ID request to its local vKMS. Since the QuSeC has already resolved and installed the required path and stores state information, the vKMS receives from the QuSeC a simple KMS identifier and forwards the request directly to the KMS (KMS_4d), which already holds the decrypted key.
@@
IV Security analysis
This section presents a theoretical security analysis of the proposed QKD trusted relay network architecture. The main objective of the analysis is to validate security properties like key confidentiality, integrity, authentication, and resilience to compromise.
To perform the validation, we first present several assumptions that mark the fundamentals of the analysis. We adopt the Dolev-Yao adversary model, where the adversary is able to intercept, modify, and change any message on any of the classical channels, either those used by the QKD algorithms, those used for key relying, or those in the control plane, although it cannot break the cryptographic primitives. We also assume that the quantum channels are secure, and that the nodes that are part of the architecture are physically secure. Apart from that, we comply with the ETSI QKD specifications, and therefore the communications between entities, namely QuSeC, vKMS, KMS, and applications, are protected through mutual authenticated and encrypted channels. In fact, the architecture is designed in compliance with the ETSI QKD 014 specification (application-to-vKMS and vKMS-to-KMS interface), the ETSI QKD 020 specification (KMS interoperability for multi-hop), and the ETSI QKD 015 specification (SDN and centralized control in QKD networks). According to the specifications, the communications must use mutually authenticated secure APIs. Additionally, in order to guarantee the confidentiality of the communication, the message exchange through REST APIs should be accordingly encrypted. These specifications and fundamentals allow the protection of communications and reduce the risk of vulnerabilities and attacks.
IV-A Architectural security
The presented architecture proposes a hierarchical design of key management. Here, there are several security layers where components are classified, consisting of the KMSs, the vKMSs, and the QuSeC. Each of the layers has different contributions and functions, which are described in this section:
  • Local KMS: Each QKD link endpoint is equipped with a local KMS, which is responsible for the key storage, reconciliation, and delivery of the secrets. The communications between the KMSs and the QKD devices are verified at link level, as each KMS is bounded to its QKD device. Higher-level processes use these KMSs for key retrieval (for applications) or key forwarding. The security features of the KMSs are the following:
@@ IV-A Architectural security
    • Link authentication: The KMS pair communicates through a mutually authenticated channel, as in the specification, ensuring protection against man-in-the-middle.
    • Isolation of keys per link: In sites with more than one link to another domain, and therefore more than one QKD device, each device has an isolated KMS, not sharing communication channel and storage.
  • Virtual KMS (vKMS): It is an abstraction of the multiple KMSs, located in an upper logical level. The vKMS manages the application queries to the QKDs in domains with more than one KMS, performing the required discovery procedure. This decision is aligned with the ETSI QKD 020 specification. Security features of vKMS:
@@
    • Policy enforcement and monitoring: The vKMS may enforce security policies within the domain toward the applications and systems consuming the cryptographic material. These policies include limitations to request rates, preventing attackers from spamming requests, or key usage policies, requesting the use of the keys from the application.
    • Interoperability across vendors: The vKMS provides an abstraction from the specific implementation of the KMSs and the manufacturers of the QKD devices toward the applications. With this layer, the application requests and communication are manufacturer-independent, reducing the risk of bugs and errors in the adaptation of each interface, supporting application mobility and avoiding vendor lock-in.
  • QuSeC: The highest level of security and management hierarchy, analogous to an SDN controller for the QKD network. It is compliant with the ETSI QKD 015, thus it has a global view of the QKD network. It is responsible for orchestration, relay path selection, and relay policy definition. Security features of QuSeC:
@@
    • No exposure of key material: The QuSeC does not retrieve, see or manipulate any QKD key material, as the applications directly access the vKMS to request these secrets. In fact, this is defined in the ETSI QKD 015 API, stating that the introduction of the SDN paradigm and a centralized controller should not imply further security risks. Even though the QuSeC is deemed secure, if it would be compromised, the separation of control and key manipulation prevents the attacker from obtaining the keys.
@@ IV-B Security properties
  • Authentication and integrity: The objective of an adversary may be to impersonate a legitimate entity or modify messages to deceive the system. In this context, the proposed architecture is designed to authenticate and protect the integrity of each critical message. This is achieved by performing mutual authentication between the entities, which is specified in the QKD specifications. In fact, an application that leverages the ETSI QKD 014 API to access a vKMS previously performs a TLS handshake, as well as between KMSs that leverage the ETSI QKD 020 API to perform a trusted relay. Thus, the authentication is robust as long as the used cryptographic systems are robust, such as certificates based on post-quantum cryptography. The integrity of the messages is also automatically provided after establishing a secure session using TLS, using sequence numbers and HMAC. Additionally, key relay is performed through OTP, which inherently provides authentication and integrity, as the encryption is one-use-only and the source is authenticated.

### v2

@@ Abstract
With the advent of quantum computing, the increasing threats to security poses a great challenge to communication networks. Recent innovations in this field resulted in promising technologies such as Quantum Key Distribution (QKD), which enables the generation of unconditionally secure keys, establishing secure communications between remote nodes. Additionally, QKD networks enable the interconnection of multi-node architectures, extending the point-to-point nature of QKD. However, due to the limitations of the current state of technology, the scalability of QKD networks remains a challenge toward feasible implementations. When it comes to long-distance implementations, trusted relay nodes partially solve the distance issue through the forwarding of the distributed keys, allowing applications that do not have a direct QKD link to securely share key material. Even though the relay procedure itself has been extensively studied, the establishment of the relaying node path still lacks a solution. This paper proposes an innovative network architecture that solves the challenges of Key Management System (KMS) identification, relay path discovery, and scalability of QKD networks by integrating Software-Defined Networking (SDN) principles, and establishing high-level virtual KMSs (vKMS) in each node and creating a new entity called the Quantum Security Controller (QuSeC). The vKMS serves the end-user key requests, managing the multiple KMSs within the node and abstracting the user from discovering the correct KMS. Additionally, based on the high-level view of the network topology and status, the QuSeC serves the path discovery requests from vKMSs, computing the end-to-end (E2E) relay path and applying security policies. The paper also provides a security analysis of the proposal, identifying the security levels of the architecture and analyzing the core networking security properties, and a performance evaluation that confirms the feasibility and scalability of the proposal.
@@ I Introduction
As quantum technologies evolve, Quantum Key Distribution (QKD) stands as a leading technology for securing communications with unconditional security [scarani2009security]. Unlike classical cryptographic mechanisms that rely on computational assumptions, QKD is based on quantum mechanics to generate cryptographic keys with information-theoretic security.
While the typical application scenario of QKD is point-to-point by nature, QKD networks interconnect multiple nodes over broader areas, enabling key delivery beyond direct optical links. However, the scalability of such networks remains a challenge, especially as the number of users, links, loops, and relay hops increase. Addressing scalability requires architectural solutions for network control and infrastructure integration strategies that reduce deployment complexity [li2025integration].
@@
Figure 1 illustrates two QKD network topologies. The linear topology (top) represents most of the state-of-the-art trusted relay deployments, built pairs of Simple QKD nodes. This approach is applicable to static end-to-end (E2E) communications in which the relay path remains constant [ribezzo2023deploying] [zhang2018large]. However, the partial mesh topology (down) illustrates a more realistic scenario, where multiple QKD nodes co-located in the same site enable the provision of dynamic relay paths.
@@
  • Path Discovery and Relay Coordination: In nodes hosting multiple KMSs, Each KMS must identify and coordinate with the correct KMS instance within the same node to perform the relay of keys.
@@
These issues stem from the limited knowledge at individual KMSs, and by the limitations of current QKD standards, which support basic functions but lack full dynamic multi-hop capabilities. To address these limitations, this paper proposes a secure multi-hop key delivery framework based on a centralized, programmable architecture leveraging Software-Defined Networking (SDN) principles. This approach abstracts complexity from applications and KMSs, introducing a global coordination mechanism to resolve paths and manage relay processes. The proposed design enables a centralized discovery and resolution of relay paths and remains compatible with existing and evolving QKD standards. Additionally, it enables seamless integration with different manufacturers and technologies.
The rest of the paper is organized as follows. Section II summarizes the background and literature research. Section III describes the proposal Section IV-A provides a security analysis and performance evaluation results of the proposed system. Section V presents the main conclusions.
@@ II Related work
A primary limitation of QKD is its restricted transmission distance typically 100–150 kilometers, due to signal attenuation and the lack of practical quantum repeaters. To overcome these constraints and support the deployment of scalable and complex network topologies, QKD networks commonly rely on trusted relay nodes, which have been widely studied [cao2022evolution] [tsai2021quantum].
@@
Many QKD deployments are based on linear topologies, where each trusted relay node hosts two QKD modules and associated KMSs, making the relay procedure relatively straightforward. For instance, [ribezzo2023deploying] presents a linear QKD network connecting cities in Italy, Slovenia, and Croatia through three links and two relays. Similarly, the Beijing-Shanghai Backbone QKD Network [zhang2018large] connects 32 nodes linearly across 2.000 kilometers. While such networks demonstrate the feasibility of large-scale deployments, their linear nature limits scalability and flexibility.
These limitations become more pronounced in practical topologies where nodes may host multiple KMSs and QKD links. In such configurations, the process of discovering the best KMSs and computing an optimal relay path becomes significantly more complex. To address this, several efforts explore distributed routing approaches for QKD networks. For example, [drif2025distributed] proposes the use of modified versions of classical routing protocols such as OSPFv2, where each node independently computes shortest path trees and maintains a local routing table to determine relay paths.
Other works pursue centralized architectures for QKD network control and key relay. For example, [sasaki2011field] proposes a Key Management Agent per node and a centralized KMS coordinating all the links and networking functions.
In this context, SDN emerges as a promising approach. By decoupling control and data planes, SDN enables centralized control, programmability, dynamic path computation, and real-time reconfiguration, features well-suited for QKD networks with varying key rates, link availability, and security constraints. Several efforts validate SDN integration in QKD infrastructures. For instance, [martin2024madqci] presents a network comprised of 28 QKD modules distributed across 9 nodes in Madrid. The network is coordinated by an SDN controller through the SDN Agents of the nodes. However, their relay nodes do not have KMS functionality for key retrieval, only extending the QKD reach. Recent studies [lopez2025enhanced] have also explored virtualized control and key management models, enhancing flexibility and interoperability through softwarized architectures.
Additionally, [bassi2023quantum] proposes an SDN-managed QKD infrastructure prototype where E2E keys are delivered via key relay procedure orchestrated by the controller, which computes suitable key relay paths. However, their description is based on a simple linear chain and the procedure of path computation is not described in detail, which limits its applicability to more complex topologies.
In line with these SDN-based approaches, the ETSI GS QKD 015 [ETSI_GS_QKD_015] specification stands as a key enabler for building interoperable and programmable QKD networks. This specification defines an interface and data models between the Software-Defined QKD nodes and the controller, formalizing the architectural principles to integrate SDN control into QKD infrastructures.
@@ III-A Architecture
  • Local KMS: KMS instance associated to a physical QKD module with standard KMS functionalities: key storage, synchronization, management and exchange. It also handles both direct key retrieval and multi-hop key relaying in accordance with instructions received from the QuSeC. From this point forward, these instances will be referred to as simply KMS.
  • QuSeC: It maintains a global view of the QKD network topology, including QKD links, node roles, and KMS associations. It responds to discovery requests from vKMS instances, computes optimal relay paths and enforces relay policies to all involved KMSs across the network. The relay path computation is performed based on SPF algorithm [dijkstra], assigning the weights of each link according to configurable criteria such as available key rate, hop count, or physical distance, allowing for flexible adaptation to different optimization goals.
The architecture is agnostic to the underlying QKD protocol and transmission medium because all entities interact through standardized REST APIs (ETSI QKD 014 and 020), ensuring interoperability across heterogeneous vendors and devices. From a scalability perspective, the design supports horizontal scalability, enabling additional QKD nodes to be integrated without compromising the network and KMS operation.
Beyond simple configuration of link weights, the QuSeC enables programmability through its northbound interface, allowing operators to program different routing or security policies dynamically. For instance, modifying the behavior of the Path Computation Element (PCE) in the QuSeC, the routing algorithm can be replaced or modified seamlessly.
An example of the proposed architecture is depicted in Figure 2. In this scenario, three nodes operate as trusted relay nodes, each hosting multiple KMS instances (each connected to a different QKD module, potentially from different vendors) to interface with separate QKD links. The fourth node acts as a simple QKD node with a single QKD module.
@@ III-B Operational flow
The proposed architecture enables E2E QKD key delivery through a mechanism that integrates centralized computation of the relay path. In addition, the mechanism are compliant with current specifications: ETSI GS QKD 014 [ETSI_GS_QKD_014] for key delivery, ETSI GS QKD 020 [ETSI_GS_QKD_020] for interoperable KMSs in horizontal key relay procedures, which is currently under development, and ETSI GS QKD 015 [ETSI_GS_QKD_015] for SDN integration. This section outlines the main operational steps from the application’s key request to the delivery of the final cryptographic key, as illustrated in Figure 2, which includes the operations required for initiating (top) and target (down) applications.
Figure 2: Overall steps of the proposed mechanism for an initiating application (top) and target application (down).
@@ III-C Message Exchange/examples
This section presents two representative message exchange sequences under different scenarios: one using a direct key delivery and the other involving a trusted relay-based relay path. These examples focus on the interaction between the applications, vKMS, KMSs, and the QuSeC during the key request and delivery process.
@@
Since no further hops are installed in the KMS, the relay process is completed. In addition, completion notifications are propagated back through the chain. Once KMS_1b receives the confirmation of successful relay completion, it responds to the initial Get Key request, returning K1 to the vKMS, which then delivers it to APP_A.
APP_B’s behavior mirrors the direct case: it sends a Get Key with ID request to its local vKMS. Since the QuSeC has already installed the required path and stores state information, the vKMS receives from the QuSeC a simple KMS identifier and forwards the request directly to the KMS (KMS_4d), which already holds the decrypted key.
@@
IV Security analysis and performance assessment
This section presents a theoretical security analysis of the proposed framework and the results of the performance evaluation.
IV-A Security analysis
To perform the validation, we first present several assumptions that mark the fundamentals of the analysis. We adopt the Dolev-Yao adversary model, where the adversary is able to intercept, modify, and change any message on the classical channels, either those used by the QKD algorithms, those used for key relying, or those in the control plane, although it cannot break the cryptographic primitives. We also assume that the quantum channels are secure, and that the nodes that are part of the architecture are physically secure. Apart from that, we comply with the ETSI QKD specifications, and therefore the communications between entities, namely QuSeC, vKMS, KMS, and applications, are protected through mutual authenticated and encrypted channels.
IV-A1 Architectural security
The presented architecture proposes a hierarchical design of key management. Here, there are several security layers where components are classified: KMSs, vKMSs, and QuSeC, with different contributions and functions:
  • Local KMS: Each QKD link endpoint is equipped with a local KMS responsible for the key storage, reconciliation, and delivery. The communications between the KMSs and the QKD devices are verified at link level, as each KMS is bounded to its QKD device. The security features of the KMSs are the following:
@@ IV-A1 Architectural security
    • Link authentication: The KMS pair communicates through a mutually authenticated channel, ensuring protection against man-in-the-middle.
    • Isolation of keys per link: In sites with more than one link, and therefore multiple QKD devices, each device has an isolated KMS, not sharing communication channel and storage.
  • Virtual KMS (vKMS): It is an abstraction of the multiple KMSs, located in an upper logical level. The vKMS manages the application queries to the QKDs, in domains with more than one KMS, performing the required KMS discovery procedure. This decision is aligned with the ETSI QKD 020 specification. Security features of vKMS:
@@
    • Policy enforcement and monitoring: The vKMS may enforce security policies within the domain toward the applications. These policies include limitations to request rates or key usage policies.
    • Interoperability across vendors: The vKMS provides an abstraction from the specific implementation of the KMSs and the manufacturers of the QKD devices toward the applications. With this layer, the application requests and communication are manufacturer-independent, reducing the risk of bugs and errors, supporting application mobility and avoiding vendor lock-in.
  • QuSeC: The highest level of security and management hierarchy, analogous to an SDN controller for the QKD network. It is compliant with the ETSI QKD 015, and it is responsible for orchestration, relay path selection, and relay policy definition. Security features of QuSeC:
@@
    • No exposure of key material: The QuSeC does not retrieve, see or manipulate any QKD key material, as the applications directly access the vKMS to request these secrets. Even though the QuSeC is deemed secure, if it would be compromised, the separation of control and key manipulation prevents the attacker from obtaining the keys.
@@ IV-A2 Security properties
  • Authentication and integrity: The objective of an adversary may be to impersonate a legitimate entity or modify messages to deceive the system. In this context, The proposed architecture is designed to authenticate and protect the integrity of each critical message. This is achieved by performing mutual authentication between the entities, which is specified in the QKD specifications. In fact, an application that leverages the ETSI QKD 014 API to access a vKMS previously performs a TLS handshake, as well as between KMSs that leverage the ETSI QKD 020 API to perform a key relay. Thus, the authentication is robust as long as the used cryptographic systems are robust, such as certificates based on post-quantum cryptography. The integrity of the messages is also automatically provided after establishing a secure session using TLS, using sequence numbers and HMAC. Additionally, key relay is performed through OTP, which inherently provides authentication and integrity, as the encryption is one-use-only and the source is authenticated.
IV-B Performance assessment
To validate the feasibility and performance of the proposed architecture, a testbed was deployed on a Kubernetes cluster hosting containerized instances of the QuSeC, vKMS, and KMS entities. Figure 5 presents the measured E2E key establishment delay for different communication types, ranging from direct path to five-hop relay scenarios. The path computation assumes equal link costs, so the routing criterion corresponds to the minimum hop count, although these weights could be dynamically adjusted to reflect metrics such as available key rate.
The results, averaged over 100 iterations per case, show a quasi-linear increase in delay with the number of hops, from about 50 ms for direct paths to around 290 ms for five hops. The most significant increase occurs between the direct and one-hop cases, as the first relay introduces both the QuSeC path-installation phase and the relay procedure itself, whereas additional hops only contribute with incremental encryption/decryption and messaging operations. These results confirm that the centralized control logic introduces negligible overhead compared with direct key exchange and that the relay mechanism scales efficiently even under multi-hop configurations.
Figure 5: Measured E2E key establishment delay.